OpenStudy (anonymous):

what is the slope of the line that passes through the given points (3 2) and (5 12)

OpenStudy (mathstudent55):

Use the slope formula, the slope of the line through points (x1, y1) and (x2, y2) is (y2 - y1)/(x2 - x1). It doesn't make a difference which point you call (x1, y1) and which you call (x2, y2). In your case, slope = (12 - 2)/(5 - 3) Simplify the fraction

OpenStudy (anonymous):

Thank you sooooo muchhhh !!!! I have another one also ! And the way you explained it , made it so easy to understand! what is an equation of the line in slope intercept form m=2/7 and the y intercept is (0 -12)

OpenStudy (mathstudent55):

The slope intercept form of an equation is y = mx + b, where m = slope (in your case m = 2/7), and the b is the y-intercept. The y-intercept is the y-coordinate of the the y-intercept ordered pair (in your case (0, 12), so b = -12) Just replace m and b with your numbers and that's it.

OpenStudy (anonymous):

Thank you !!